OpenStudy (anonymous):

The Burlington Phone Company charges 8 cents per minute for any phone call within the continental US. Its rival, the Try Our Phone Instead Company, charges only 4 cents per minute for any phone call within the continental US, but it also charges a 28-cent "per call charge" every time you make a call. Write a system of equations representing these facts. Calculate how long a phone call must be to make the phone charges of both companies identical. If your average phone call is 11 minutes long, which company should you choose? Make sure you show the basis for your decision.

OpenStudy (anonymous):

assume both calls are x minutes long. Charge for Burlington, 8x charge for Try Our Phone is 28+4x they should be equal so 8x=28+4x x=7 Any call longer than 7 minutes then Try Our Phone will work out cheaper

OpenStudy (anonymous):

thanks